# JEL: C81, O32

### v3

@@ 2.1. Context
Declining pharmaceutical research productivity has been a topic of policy concern since the late 1990s. In the past thirty years, discussions have largely centered on two sets of empirical patterns, which Cockburn, (2006) summarizes as follows: “More and more money is being invested in R&D, but the rate at which new drugs are introduced is failing to keep pace.” R&D, in this setting, refers to a particular type of biomedical research: investments in preclinical and, especially, clinical studies intended to translate discoveries made in basic science into commercialized medicines suitable for use in human patients. That inputs per new drug appear to be rising sharply over time is viewed, generally, as an established fact in the empirical productivity literature (Bloom et al.,, 2020; Goldin et al.,, 2024).
@@
Measures of productivity in this setting are somewhat crude. The measure of outputs in the most influential articulation of this phenomenon, by Scannell et al., (2012), is the number of “new molecular entities”—innovative drugs—approved by the U.S. Food and Drug Administration. [^11] Technically, a new molecular entity is a drug that contains no previously approved active moiety. In practice, this means that every active element of the drug (e.g., acetaminophen and caffeine in the commonly used over-the-counter migraine medicine Excedrin) must be novel. As Bloom et al., (2020) note, that this count has been stagnant or declining is viewed as an established fact. [^12] In data on new molecular entities approved by the FDA between 2005 and 2021, we observe a small increase in the number of approvals, largely driven by an uptick in approvals in 2020 and 2021. In earlier periods of data, we confirm that there is no discernible trend in new molecular entity approvals.
Measures of inputs are, typically, some estimate of R&D expenditures. The simplest construction of a measure of R&D expenditure scales the (inflation-adjusted) cost of a clinical trial by the number of clinical trials. Recent evidence on these objects comes from Bloom et al., (2020) and Sertkaya et al., (2024). Bloom et al., (2020) find a sharp increase in the number of clinical trials, as measured in publication data over time. Using granular data on clinical trial site contracts and internal records of the U.S. Food and Drug Administration, Sertkaya et al., (2024) find that the cost of clinical trials and the cost of drug development was roughly constant between 2000 and 2018.
@@ 2.2.1. Trial Registries and Proprietary Databases
Nearly all clinical trials are conducted under the auspices of a regulator that requires both pre-registration and results disclosure. In the United States, the Food and Drug Administration Amendments Act (FDAAA), as passed in 2007 and revised in a ”Final Rule” that took effect in 2017, imposes these requirements on sponsors of regulated trials. Nominally, effected sponsors are required to register their studies on “ClinicalTrials.gov.” For nearly all studies, sponsors must also report the results of a trial within a designated period after completion. Moreover, increasingly, medical journals require registration as a pre-condition of publication (Laine et al.,, 2007).
@@
In practice, there is widespread noncompliance with these regulations and requirements. Deborah Zarin, then-director of ClinicalTrials.gov, notes that “funders, sponsors, and institutional review boards continue to allow unregistered trials or trials with late registration to be conducted, and some journals continue to allow the results of such trials to potentially be published” (Zarin et al.,, 2017). Noncompliance with federal reporting requirements is similarly widespread. In a series of papers assessing compliance over time DeVito et al., (2020) and DeVito and Goldacre, (2021) find that roughly 60 percent of trials registered on ClinicalTrials.gov fail to report required data elements at the time of their reporting deadline. [^13] Note that noncompliance in this industry persists despite the existence of high statutory penalties. Sponsors can be fined up to \$10,000 per day for failing to register studies and disclose results. See the FDA Amendments Act of 2007 (FDAAA) and its Final Rule issued in 2021, at 42 U.S.C. § 282(j)(5)(C)(i). To date, no fines have been collected, but the FDA now sends notices of noncompliance to delinquent sponsors (Stephenson,, 2021). Notices are available here: https://www.fda.gov/science-research/fdas-role-clinicaltrialsgov-information/clinicaltrials gov-notices-noncompliance-and-civil-money-penalty-actions
@@
From the perspective of our study, there are three issues with the use of registry data. First, requirements for study registration have changed over time. Table B.1 lists five changes between 2000 and 2020 that altered the set of studies that should appear in a complete version of the database. Second, sponsors engage in considerable “back-filling” of past studies. Zarin et al., (2017) find that a large proportion of trials are registered after the trial start date. [^14] Often back-filling relates to changes in organizational disclosure policy. In a version of ClinicalTrials.gov current through 31 May 2023, Boehringer Ingelheim—the largest privately held pharmaceutical company in the world—appears to have posted 804 studies in 2014, only 72 of which were started in 2014. In 2013, Boehringer registered 101 studies. In the fifteen preceding years, since the creation of the database, Boehringer posted 872 records in total. We selected this case study based on a discussion in Zarin et al., (2017). Third, not all records in ClinicalTrials.gov and similar registries are, in fact, clinical trials (Tse et al.,, 2018). As a result, it is difficult to determine whether changes in the number of records reflect actual growth in scientific research or a change in the propensity to register studies (trials and non-trial studies).
@@
In Appendix B, we illustrate the difficulty of recovering accurate measures of trends in the quantity of clinical trials from this database. We show that small changes in data construction yield meaningfully different estimates of trial levels and that many records are not in fact clinical trials. Moreover, due to widespread data gaps resulting from noncompliance in reporting of results and study attributes, there is little reliable information about characteristics of registered studies.
@@
Other databases, such as the PDQ Cancer Information database maintained by the U.S. National Cancer Institute and used in papers including Budish et al., (2015), are disease-specific and, thus, insufficiently general for our purposes. A small number of firms (notably, GlaxoSmithKline) have disclosed all clinical trials over a longer period of time. [^15] See GlaxoSmithKline’s registry here: https://www.gsk-studyregister.com/ Of course, records from a single firm are not informative about aggregate trends, especially because many investigational drugs change hands frequently during the course of development (Cunningham et al.,, 2021).
@@
Researchers, instead, often draw on proprietary databases. Although these databases are sufficient for certain research questions—for example, related to changes in the ownership of investigational drugs (Cunningham et al.,, 2021, using Pharmaprojects data) or on the link between insurance policy and drug development (Agha et al.,, 2022, using Cortellis data)—they are unsuitable for our particular productivity measurement exercise. Our aim is to construct a census that allows for comparison of quantities across points in time. First, we are unable to determine, with specificity, how records are identified and catalogued based on available documentation. Second, these databases take public registries as inputs and, thus, risk adopting the same shifts in reporting over time (see e.g., the list of registry inputs for one frequently-used database, Cortellis,, 2024).
@@ 2.2.2. Publication Data
Publication data provide a natural alternative. [^16] In Section 4, we discuss the two natural concerns with measures constructed from publication data—publication bias and shifts in reporting norms over time. Neither is likely to affect the interpretation of trends in our data. The database PubMed / MEDLINE indexes roughly $34$ million publications, a near universe of published biomedical research. [^17] See Section A.1 for further details on on our treatment of the PubMed data. The challenge is to classify the specific forms of research captured by this corpus. Of these $34$ million records, which are clinical trials?
The existing literature takes several approaches. In health sciences—where data on the universe of clinical trials are relevant for the construction of meta-analyses and the development of prescribing guidelines—researchers typically employ over-inclusive methods with high true positive rates. To use one prominent example, Cochrane, a British organization that synthesizes the findings of medical research, developed the “Cochrane RCT Classifier,” a machine learning-based classifier for retrieving randomized controlled trials. The Cochrane approach successfully identifies trials with a 99 percent true positive rate. For every eight true positives, however, it returns 92 false positives. See Thomas et al., (2021) for details.
@@ 2.3. Objective
As a baseline, we assess the performance of several standard machine learning algorithms. The results are displayed in Figure 2.2. Further details on the construction of this figure are given in Section C.2. Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits. Performance is measured in the testing sample split. For feature vectors, we use either TF-IDF embeddings computed in the corpus of abstracts in the hand-labeled sample or the embeddings of each abstract obtained from the SentenceTransformer language model (Reimers and Gurevych,, 2019). We find that the performance of the standard machine learning algorithms is again unsuitable for our application. At a 90% true positive rate, the best performing model identifies 50 true positives for every 50 false positives.
@@ 2.4.1. Prompt Design
This exercise is instructive on three margins. First, it highlights the highest performing prompts. Second, it indicates particular types of errors in categorization—which suggest opportunities for more precise language in a prompt. Third, it draws attention to differences in the performance of the two models. Observe that, in the “Other” error type, we include a sub-type called “overly literal interpretation of inclusion criteria.” We primary record errors of this type for Prompt 3, which asked the model to return a completion that described why, or why not, a record was classified as being in our sample. Here, GPT-3.5 makes 30 such errors—classifying records as FALSE by recapitulating the sample definition provided. GPT-4 makes two errors. The insight that emerges from this exercise, then, is that different prompt structures may be preferable depending on the model used, and that, by and large, GPT-4 is better suited for tasks that require (the resemblance of) more sophisticated reasoning.
@@ 2.4.2. Fine-Tuning
BigBird uses a sparse attention mechanism to reduce the computational cost of processing long text sequences. [^23] The computational complexity of regular attention is quadratic in sequence length, while BigBird reduces this to linear complexity. This enables efficient handling of long documents up to 4,096 tokens. In contrast, standard BERT models can only process 512 tokens. BigBird is useful because abstracts of clinical trials regularly exceed 512 tokens. Prior to pre-training, the model was warm-started from the RoBERTa checkpoint (Liu et al.,, 2019) and then pre-trained on the standard BERT corpus.242424 The standard BERT corpus consists of the Books (Zhu et al.,, 2015), CC-News (Guu et al.,, 2020), Stories (Trinh and Le,, 2019), and Wikipedia (Wikimedia Foundation,, 2023) datasets. We evaluate two model sizes: BigBird Base (125 million parameters) and BigBird Large (355 million parameters).
@@ 2.5. Performance
We choose three thresholds according to the stringency with which they enforce the sample restrictions. The estimated true and false positive rates associated with these points are displayed in Figure 2.4 and labeled as “Conservative,” “Moderate,” and “Liberal.” Our preferred sample is associated with the conservative threshold. [^26] The liberal model may be particularly useful for conducting literature reviews, where a near-perfect true positive rate is needed. For every 82 true positives, the conservative model identifies 18 false positives. [^27] In the test data, the conservative model assigns incorrect labels to 27 of 993 papers. We conduct an error analysis. See Section C.4 for further details. In 13 cases, there is a clear error. In 14 cases, however, errors are associated with records that were difficult to categorize for a human labeller. For example, PubMed record 32737793 is flagged as satisfying 2.1 with all three thresholds, but is a literature review. By contrast, PubMed record 27880726 was categorized as satisfying 2.1 twice by a human labeller. It is excluded from both the conservative and moderate model-generated samples. On inspection, the abstract does not explicitly state that the study enrolled human subjects, but hints that it may have been conducted in an animal model. Review of the associated full text confirms that this study did, in fact, enroll only rats. We report results associated with the moderate and liberal thresholds in the appendix as tests of robustness. Our final, conservative, sample consists of 152,027 publications classified as satisfying the sample restrictions.

### v5

@@ 2.1. Context
Declining pharmaceutical research productivity has been a topic of policy concern since the late 1990s. In the past thirty years, discussions have largely centered on two sets of empirical patterns, which Cockburn, (2006) summarizes as follows: “More and more money is being invested in R&D, but the rate at which new drugs are introduced is failing to keep pace.” In these exercises, the empirical object of interest is a measure of investments in preclinical and, especially, clinical studies intended to translate discoveries made in basic science into commercialized medicines suitable for use in human patients.
The “puzzle” as characterized by Cockburn, (2006) and others centers on the apparent decline in the efficiency of pharmaceutical firms tasked with translational research. Hypotheses for the decline are wide-ranging. Ruffolo, (2006) emphasizes shifts in managerial incentives. Cockburn, (2006) highlights “vertical disintegration” of the pharmaceutical industry. Scannell et al., (2012) points to rising regulatory thresholds. Myers and Pauly, (2019) and Bloom et al., (2020) suggest that the “low-hanging fruit”—high-efficacy, easy-to-test drug candidates—have already been brought to market. Though the candidate mechanisms in these discussions differ, declining productivity in the pharmaceutical sector is treated as a fact (Bloom et al.,, 2020; Goldin et al.,, 2024). [^11] Scannell et al., (2012) coin the term “Eroom’s Law”—Moore’s Law in computing, spelled backwards—to described this trend. Policy analyses and academic surveys (e.g., Goldin et al., (2024)) have characterized both Eroom’s Law and Moore’s Law as well-supported stylized facts.
@@
The measures of industry productivity central to these debates are somewhat crude. For the pharmaceutical industry, a standard measure of output is the number of “new molecular entities”—innovative drugs—approved each year by the U.S. Food and Drug Administration. [^12] Technically, a new molecular entity is a drug that contains no previously approved active moiety. In practice, this means that every active element of the drug (e.g., acetaminophen and caffeine in the commonly used over-the-counter migraine medicine Excedrin) must be novel. This is the same output measure used in discussions of changes in federal research policy. See, for example, Harris, (2011). As Bloom et al., (2020) note, that this count has been stagnant or declining over time is viewed as an established fact. [^13] In data on new molecular entities approved by the FDA between 2005 and 2021, we observe a small increase in the number of approvals, largely driven by an uptick in approvals in 2020 and 2021. In earlier periods of data, we confirm that there is no discernible trend in new molecular entity approvals.
Measures of inputs are, typically, some estimate of R&D expenditures. The most widely cited analysis of pharmaceutical research productivity, Scannell et al., (2012), uses a measure of annual, aggregated industry-wide R&D expenditures, published by the trade organization Pharmaceutical Research and Manufacturers of America (PhRMA). [^14] These PhRMA estimates are used more widely. See e.g., Myers and Pauly, (2019) for use in an economics paper on pharmaceutical productivity. See e.g., Harris, (2011) for a discussion in the popular press. There are few other sources of R&D expenditure data for the pharmaceutical industry. DiMasi et al., (2003) and DiMasi et al., (2016) provide, perhaps, the most commonly cited estimates of drug development costs, derived from internal records of a select set of firms. In an editorial, Frank, (2003) describes these estimates as ”a matter of heated debate since they were first made public.” [^15] Light and Warburton, (2005), for example, criticizes these estimates derived from industry data as ”nonrandom and small” and characterizes the findings as based on ”unverifiable industry data.”
Given these substantial limitations of industry-wide aggregate estimates, a natural alternative is to decompose aggregate R&D into its component parts. The simplest such decomposition would develop estimates of the cost of pre-clinical and clinical studies and scale this quantity by the number of such clinical trials. In practice, nearly all work focuses on clinical trials and assumes that pre-clinical trials contribute a fixed proportion of total development costs (DiMasi et al.,, 2003, 2016; Sertkaya et al.,, 2024). [^16] Ideally, estimates of clinical trial costs would be inflation-adjusted. To date, there are no price indices that are appropriate for such an adjustment. See Berndt and Cockburn, (2013) for an effort to construct such a price index. See Sertkaya et al., (2016) for a discussion of efforts to construct cost estimates, including limitations associated with this type of adjustment. Two recent estimates of these objects come from Bloom et al., (2020) and Sertkaya et al., (2024). In an influential analysis, Bloom et al., (2020) document a sharp increase in the number of clinical trials over time, using data constructed using scientific publications. In Bloom et al., (2020) and in a recent survey article, Goldin et al., (2024), this fact is central to the argument that pharmaceutical productivity is declining. Using granular data on clinical trial site contracts and internal records of the U.S. Food and Drug Administration, Sertkaya et al., (2024) find both that the cost of clinical trials and the cost of drug development were roughly constant between 2000 and 2018.171717The data used in the Sertkaya et al., (2024) analysis have previously been used by a small set of other researchers—notably, Azoulay, (2004) and Berndt and Cockburn, (2013). As of 2019, the firm that collects data on clinical trial contracts has been acquired. As of 2022, the data are no longer available for research use. Correspondence on file with the authors.
@@ 2.2.1. Trial Registries and Proprietary Databases
Nearly all clinical trials are conducted under the auspices of a regulator that requires both pre-registration and results disclosure. In the United States, the Food and Drug Administration Amendments Act (FDAAA), as passed in 2007 and revised in a ”Final Rule” that took effect in 2017, imposes these requirements on sponsors of regulated trials. Technically, affected sponsors are required to register their studies on “ClinicalTrials.gov.” For nearly all studies, sponsors must also report the results of a trial within a designated period after completion. Moreover, increasingly, medical journals require registration as a pre-condition of publication (Laine et al.,, 2007).
@@
In practice, there is widespread noncompliance with these regulations and requirements—by private firms, academic institutions, and the federal government itself. [^18] There have been many extremely careful, detailed analyses of ClinicalTrials.gov non-compliance. See Anderson et al., (2015) for one comprehensive analysis. The FDAAA Trials Tracker—https://fdaaa.trialstracker.net/—provides up-to-date, record-by-record details on non-compliance. Deborah Zarin, then-director of ClinicalTrials.gov, notes that “funders, sponsors, and institutional review boards continue to allow unregistered trials or trials with late registration to be conducted, and some journals continue to allow the results of such trials to potentially be published” (Zarin et al.,, 2017). Noncompliance with federal reporting requirements is similarly widespread. In a series of papers assessing compliance over time, DeVito et al., (2020) and DeVito and Goldacre, (2021) find that roughly 60 percent of trials registered on ClinicalTrials.gov fail to report required data elements at the time of their reporting deadline. [^19] Note that noncompliance in this industry persists despite the existence of high statutory penalties. Sponsors can be fined up to \$10,000 per day for failing to register studies and disclose results. See the FDA Amendments Act of 2007 (FDAAA) and its Final Rule issued in 2021, at 42 U.S.C. § 282(j)(5)(C)(i). To date, no fines have been collected, but the FDA now sends notices of noncompliance to delinquent sponsors (Stephenson,, 2021). Notices are available here: https://www.fda.gov/science-research/fdas-role-clinicaltrialsgov-information/clinicaltrials gov-notices-noncompliance-and-civil-money-penalty-actions
@@
From the perspective of our study, there are three issues with the use of registry data. First, requirements for study registration have changed over time. Table B.1 lists five changes between 2000 and 2020 that altered the set of studies that should appear in a complete version of the database. Oostrom, (2024) documents, in a case study of psychiatric drugs, large changes in reporting propensity over this time period. [^20] More details on the Oostrom, (2024) comparison of registry data and publication data appear in Section 4. Second, sponsors engage in considerable “back-filling” of past studies. Zarin et al., (2017) find that a large proportion of trials are registered after the trial start date. [^21] Often back-filling relates to changes in organizational disclosure policy. In a version of ClinicalTrials.gov current through 31 May 2023, Boehringer Ingelheim—the largest privately held pharmaceutical company in the world—appears to have posted 804 studies in 2014, only 72 of which were started in 2014. In 2013, Boehringer registered 101 studies. In the fifteen preceding years, since the creation of the database, Boehringer posted 872 records in total. We selected this case study based on a discussion in Zarin et al., (2017). Third, not all records in ClinicalTrials.gov and similar registries are, in fact, clinical trials (Tse et al.,, 2018). As a result, it is difficult to determine whether changes in the number of records reflect actual growth in scientific research or a change in the propensity to register studies (trials and non-trial studies).
@@
In Appendix B, we illustrate the difficulty of recovering accurate measures of trends in the quantity of clinical trials from this database. In particular, we show that small changes in data construction yield meaningfully different estimates of trial levels and that many records are not in fact clinical trials. Moreover, due to widespread data gaps resulting from noncompliance in reporting of results and study attributes, there is little complete information on characteristics of registered studies. [^22] See Alsan et al., (2024) for a discussion of gaps in the reporting of the racial composition of studies.
@@
Other databases, such as the PDQ Cancer Information database maintained by the U.S. National Cancer Institute and used in papers including Budish et al., (2015), are disease-specific and thus insufficiently general for our purposes. A small number of firms (for example, GlaxoSmithKline) have disclosed all clinical trials over a longer period of time. [^23] See GlaxoSmithKline’s registry here: https://www.gsk-studyregister.com/. Gibson, (2004) provides an overview of the lawsuit that led to the development of this registry. Of course, records from a single firm are not informative about aggregate trends, especially because many investigational drugs change hands frequently during the course of development (Cunningham et al.,, 2021).
@@
Researchers, instead, often draw on proprietary databases. Although these databases are extremely useful for certain research questions—for example, regarding changes in the ownership of investigational drugs (Cunningham et al.,, 2021, using Pharmaprojects data) or on the link between insurance policy and drug development (Agha et al.,, 2022, using Cortellis data)—they are unsuitable for analyses of either research trends (as in this paper) or summaries of the evidence base relevant for clinical decision-making (as in the health services literature). Our aim is to construct a census that allows for comparison of quantities across points in time. First, we are unable to determine, with specificity, how records are identified and cataloged based on available documentation. Second, these databases take public registries as inputs and, thus, risk adopting the same shifts in reporting over time (see e.g., the list of registry inputs for one frequently-used database, Cortellis,, 2024).
@@ 2.2.2. Publication Data
Publication data provide a natural alternative. [^24] In Section 4, we discuss the two natural concerns with measures constructed from publication data—publication bias and shifts in reporting norms over time. Our data, a recent analysis by Oostrom, (2024), and the relevant institutional context suggest, together, that neither is likely to threaten our interpretation of trends in publication data. The database PubMed / MEDLINE indexes roughly $34$ million publications, a near universe of published biomedical research. [^25] See Section A.1 for further details on our treatment of PubMed data. The challenge is to classify the specific forms of research captured by this corpus. Of these $34$ million records, which are clinical trials?
The existing literature takes several approaches. In health sciences—where data on the universe of clinical trials are relevant for the construction of meta-analyses and the development of prescribing guidelines—researchers typically employ overly-inclusive methods characterized by high true positive rates. To use one prominent example, Cochrane, a British organization that synthesizes the findings of medical research, developed the “Cochrane RCT Classifier,” a machine learning-based classifier for retrieving randomized controlled trials. The Cochrane approach successfully identifies trials with a 99 percent true positive rate. For every eight true positives, however, it returns 92 false positives. Thomas et al., (2021) provides details on the Cochrane approach.
@@ 2.3. Objective
As a baseline, we assess the performance of several standard machine learning algorithms. The results are displayed in Figure 2.2. Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits. Performance is measured in the testing sample split. For feature vectors, we use either TF-IDF embeddings computed in the corpus of abstracts in the hand-labeled sample or the embeddings of each abstract obtained from the SentenceTransformer language model (Reimers and Gurevych,, 2019). We find that the performance of the standard machine learning algorithms is again unsuitable for our application. At a 90% true positive rate, the best performing model identifies 50 true positives for every 50 false positives.
@@ 2.4.1. Prompt Design
This exercise is instructive on three margins. First, it highlights the highest performing prompts. Second, it indicates particular types of errors in categorization—which suggest opportunities for more precise language in a prompt. Third, it draws attention to differences in the performance of the two models. Observe that, in the “Other” error type, we include a sub-type called “overly literal interpretation of inclusion criteria.” We primarily record errors of this type for Prompt 3, which asked the model to return a completion that described why, or why not, a record was classified as being in our sample. Here, GPT-3.5 makes 30 such errors—classifying records as FALSE by recapitulating the sample definition provided. GPT-4 makes two errors. The insight that emerges from this exercise, then, is that different prompt structures may be preferable depending on the model used, and that, by and large, GPT-4 is better suited for tasks that require (the resemblance of) more sophisticated reasoning.
@@ 2.4.2. Fine-Tuning
BigBird uses a sparse attention mechanism to reduce the computational cost of processing long text sequences. [^31] The computational complexity of regular attention is quadratic in sequence length, while BigBird reduces this to linear complexity. This enables efficient handling of long documents up to 4,096 tokens. In contrast, standard BERT models can only process 512 tokens. BigBird is useful because abstracts of clinical trials regularly exceed 512 tokens. Prior to pre-training, the model was warm-started from the RoBERTa checkpoint (Liu et al.,, 2019) and then pre-trained on the standard BERT corpus.323232 The standard BERT corpus consists of the Books (Zhu et al.,, 2015), CC-News (Guu et al.,, 2020), Stories (Trinh and Le,, 2019), and Wikipedia (Wikimedia Foundation,, 2023) datasets. We evaluate two model sizes: BigBird Base (125 million parameters) and BigBird Large (355 million parameters).
@@ 2.5. Performance
We choose three thresholds according to the stringency with which they enforce the sample restrictions. The estimated true and false positive rates associated with these points are displayed in Figure 2.4 and labeled as “Conservative,” “Moderate,” and “Liberal.” Our preferred sample is associated with the conservative threshold. [^34] The liberal model may be particularly useful for conducting literature reviews, where a near-perfect true positive rate is needed. For every 82 true positives, the conservative model identifies 18 false positives. [^35] In the test data, the conservative model assigns incorrect labels to 27 of 993 papers. We conduct an error analysis. See Section C.3 for further details. In 13 cases, there is a clear error. In 14 cases, however, errors are associated with records that were difficult to categorize for a human labeller. For example, PubMed record 32737793 is flagged as satisfying 2.1 with all three thresholds, but is a literature review. By contrast, PubMed record 27880726 was categorized as satisfying 2.1 twice by a human labeller. It is excluded from both the conservative and moderate model-generated samples. On inspection, the abstract does not explicitly state that the study enrolled human subjects, but hints that it may have been conducted in an animal model. Review of the associated full text confirms that this study did, in fact, enroll only rats. All results that follow are robust to the use of the moderate and liberal thresholds. Our final, conservative, sample consists of 152,027 publications classified as satisfying the sample restrictions.